# A Review of Nine Physics Engines for Reinforcement Learning Research

### v1

@@ III-A Popularity Analysis
Figure 2 shows, that the usage of Nvidia Isaac has seen substantial growth from 2021 to 2022, with mentions increasing from 20 to 144. Mujoco maintained its popularity from 2020 to 2021, with mentions remaining high at 695 and 709, respectively. However, there was a significant drop in usage in 2022 to 761 mentions and further to 458 mentions in 2023. Chrono’s usage has remained relatively stable over the years, with only minor fluctuations in mentions from 2016 to 2023. Gazebo experienced fluctuations in usage, with a peak of 369 mentions in 2020 and a subsequent decrease to 356 mentions in 2022. In 2023, there was a noticeable increase to 199 mentions. Webots had a consistent number of mentions from 2011 to 2020, with 72 mentions in 2020. However, in recent years, there has been a decline in usage, with 36 mentions in 2023. ML Agents saw a significant increase in mentions from 2019 to 2022, with a peak of 163 mentions. In 2023, there was a decrease to 68 mentions. Brax’s usage has experienced fluctuations, with a peak of 76 mentions in 2022, followed by a decrease to 62 mentions in 2023. These year-to-year differences in usage reflect the evolving preferences and trends within the robotics research community.
@@ VI-A Author contributions
This research was completed within the scope of MicrocosmAI, which made this project possible. For further information and access to the code, please visit microcosm.ai. M.K. made the main writing contribution and organized the research and writing process. C.W. contributed to the methodology, crawling algorithm and popularity comparison. H.H. made contributions to the Chrono, Gazebo, ODE and Webots chapters. J.M., E.B. and the larger MicrocosmAI research project contributed expertise, feedback and a framework for supervision. All authors researched data and literature and contributed substantially to the conceptualization of the submitted version.

### v2

@@ III-A Popularity Analysis
Figure 2 shows, that the usage of Nvidia Isaac has seen substantial growth from 2021 to 2023, with mentions increasing from 20 to 286. Mujoco maintained its popularity from 2020 to 2023, with mentions remaining high at 695 and 880, respectively. Chrono’s usage has remained relatively small in comparison but stable over the years, with only minor fluctuations in mentions from 2016 to 2023. Gazebo experienced fluctuations in usage, but overall retained a stable number of citations between 374 in 2023 and 324 in 2019. Webots had a consistent number of mentions from 2011 to 2023, with 72 mentions in 2023. ML Agents saw a significant increase in mentions from 2019 to 2022, with a peak of 163 mentions. However in 2023, there was a slight decrease to 145 mentions. Brax’s usage has steadily been increasing over recent years to 107 mentions 2023 since its publication in 2021. These year-to-year differences in usage reflect the evolving preferences and trends within the robotics and RL research communities.
@@ VI-A Author contributions
This research was completed within the scope of the MicrocosmAI project [^27] https://microcosm.ai/, which made this project possible. M.K. made the main writing contribution and organized the research and writing process. C.W. contributed to the methodology, crawling algorithm and popularity comparison. H.H. made contributions to the Chrono, Gazebo, ODE and Webots chapters. J.M., E.B. and the larger MicrocosmAI research project contributed expertise, feedback and a framework for supervision. All authors researched data and literature and contributed substantially to the conceptualization of the submitted version.
VII Acknowledgements
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) — 456666331.